# A Gift from the Integration of Discriminative and Diffusion-based Generative Learning: Boundary Refinement Remote Sensing Semantic Segmentation

### v1

@@ Abstract
Remote sensing semantic segmentation must address both what the ground objects are within an image and where they are located. Consequently, segmentation models must ensure not only the semantic correctness of large-scale patches (low-frequency information) but also the precise localization of boundaries between patches (high-frequency information). However, most existing approaches rely heavily on discriminative learning, which excels at capturing low-frequency features, while overlooking its inherent limitations in learning high-frequency features for semantic segmentation. Recent studies have revealed that diffusion generative models excel at generating high-frequency details. Our theoretical analysis confirms that the diffusion denoising process significantly enhances the model’s ability to learn high-frequency features; however, we also observe that these models exhibit insufficient semantic inference for low-frequency features when guided solely by the original image. Therefore, we integrate the strengths of both discriminative and generative learning, proposing the Integration of Discriminative and diffusion-based Generative learning for Boundary Refinement (IDGBR) framework. The framework first generates a coarse segmentation map using a discriminative backbone model. This map and the original image are fed into a conditioning guidance network to jointly learn a guidance representation subsequently leveraged by an iterative denoising diffusion process refining the coarse segmentation. Extensive experiments across five remote sensing semantic segmentation datasets (binary and multi-class segmentation) confirm our framework’s capability of consistent boundary refinement for coarse results from diverse discriminative architectures. The source code will be available at https://github.com/KeyanHu-git/IDGBR.
Index Terms:
Remote sensing semantic segmentation, boundary refinement, diffusion model, discriminative model, coarse segmentation refinement
{textblock*}
6cm(-1.5cm,-6.3cm) This manuscript is currently under review.
@@ I Introduction
Semantic segmentation aims to classify each pixel in an image, thereby partitioning the image into multiple regions with specific semantic meanings. In remote sensing, semantic segmentation is also known as land use or land cover classification, representing a fundamental task in remote sensing image analysis. Semantic segmentation of remote sensing images provides crucial support for numerous remote sensing applications, including geographic information updates, environmental monitoring, urban planning, among others.
@@
In recent years, with the development of deep learning technology, semantic segmentation methods utilizing deep learning for remote sensing images have overcome the performance bottlenecks of traditional pixel-oriented or object-oriented segmentation approaches [1][2][3]. Although these methods have significantly improved segmentation performance, they often fail to produce satisfactory boundary segmentation results [4]. Most studies attribute the difficulty of boundary region segmentation to the fact that segmentation models rely on large-scale features to capture contextual information from spatial neighborhoods [5][6]. Compared to local features, large-scale features indeed provide richer semantic information. However, large-scale features are typically obtained from low-resolution feature maps, large receptive fields, long-range pixel relationships, or global pixel associations, thereby sacrificing local detail information. To address this issue, many studies have attempted to combine contextual semantic feature learning with boundary-sensitive local feature learning. Bertasius et al. [7] enhanced the model’s boundary localization awareness by jointly optimizing pixel-wise affinity loss and semantic segmentation loss. Ding et al. [8] introduced an additional boundary category to be optimized along with semantic categories, and utilized a directed acyclic graph structure to use the learned boundaries to isolate contextual information propagation between different regions. Wang et al. [9] designed a boundary-guided multi-scale feature fusion module to enhance boundary information preservation in multi-scale features. Considering the variability of boundary pixels between different categories, Wu et al. [10] pulled each boundary pixel closer to its nearest local category center while pushing it away from surrounding categories.
@@
However, existing studies primarily focus on enhancing intra-class similarity and inter-class separability in boundary regions under discriminative classification strategies, without adequately addressing the negative effects of discriminative learning on boundary segmentation. Boundaries represent local transitions between adjacent regions in the spatial domain. From a signal frequency perspective, boundaries correspond to high-frequency components in the frequency domain. Studies have shown that Convolutional Neural Networks (CNNs) can capture high-frequency components in image features [11]. However, discriminative classification tends to make CNNs prioritize learning low-frequency components while suppressing high-frequency ones [12]. This leads to insufficient feature learning for boundary pixels, thereby increasing the likelihood of misclassification. From a probability theory perspective, discriminative classification directly learns the posterior distribution, disregarding the inherent characteristics of the data itself and focusing solely on minimizing overall classification errors in the pixel space for given input data [13]. Due to the small proportion of boundary pixels in the pixel space, errors in boundary pixel classification are often overlooked or suppressed during the discriminative learning process [14].
@@
Unlike discriminative learning, generative learning aims to produce high-fidelity samples by modeling the joint distribution of input data [13]. Therefore, theoretically, generative learning needs to accurately model both low-frequency content structures and high-frequency edge details in images, without neglecting or suppressing the learning of high-frequency components as discriminative learning does. However, for generative models such as VAEs and GANs, learning the distribution of high-frequency components remains challenging [15]. In recent years, diffusion probabilistic models have achieved breakthrough progress in generative tasks by gradually adding noise to data and learning the reverse denoising process, thereby generating high-quality samples [16][17][18]. In image generation tasks, studies have found that diffusion generative models enhance the learning of high-frequency components in images [19]. This paper also observes similar experimental phenomena in semantic segmentation tasks: by using remote sensing images as conditional information and applying diffusion noise training with Stable Diffusion [16] to segmentation label map encoded features, more realistic segmentation boundaries can be obtained. As shown in Fig. 1 (a), when we transform the segmentation maps generated by diffusion models into the frequency domain via Fourier transform, compared to discriminative models, diffusion generative models produce segmentation results in boundary regions that appear closer to the geometric contours of targets. However, from the perspective of overall segmentation metrics, there remains a significant gap in semantic reasoning accuracy between the segmentation results of diffusion generative models and discriminative models [20], as shown in Fig. 1 (b).
Figure 1: Comparison between the diffusion generative model and the discriminative model on remote sensing image semantic segmentation tasks. (a) From the visualization results, the diffusion generative model produces segmentations that more closely adhere to the target boundaries; (b) From overall segmentation metrics, there exists a significant gap in semantic reasoning accuracy between the two model types.
@@
In summary, discriminative learning excels at capturing low-frequency large-scale semantic features but struggles with accurately learn high-frequency boundary features; conversely, diffusion generative learning can effectively capture high-frequency boundary features but lacks accuracy in inferring semantic features. Based on the above analysis and observations, we propose the Integration of Discriminative and diffusion-based Generative learning for Boundary Refinement (IDGBR) framework, which includes a diffusion generative model, a discriminative model, and a conditional guidance network. First, to ensure semantic accuracy of the overall segmentation, IDGBR uses coarse segmentation results predicted by a discriminative model as semantic guidance conditions for diffusion generative learning. Although these coarse segmentation results have inaccurate boundary regions, they provide reliable semantic information for interior regions. Second, IDGBR learns joint guidance representations from coarse segmentation results and remote sensing images through a conditional guidance network, which are input to the diffusion generative model via residual connections, promoting more accurate inference of semantic features and boundary features by the generative model. Finally, we also use a simple regularization technique that explicitly aligns diffusion generative representations with joint guidance representations at intermediate network layers, greatly improving the training efficiency of diffusion generative semantic segmentation.
The main contributions of this paper are summarized as follows:
@@
  • We introduce a regularization technique to explicitly align diffusion generative representations with joint guidance representations, improving the training efficiency of diffusion generative semantic segmentation.
@@ II-A General Semantic Segmentation
Remote sensing semantic segmentation often faces challenges such as insufficient boundary refinement [21],[22],[23],[24], semantic ambiguity [25], and missed segmentation of small-scale targets [6],[26],[27] due to large intra-class differences, low inter-class discrimination, and class imbalance in remote sensing features, as well as the large size, significant scale variations, and high scene complexity of remote sensing images. To address these issues, researchers have conducted a series of explorations.
CNNs represent the classical architecture in remote sensing semantic segmentation, with numerous research efforts building upon their strong local feature extraction capabilities to enhance the perception of multi-scale contextual information and boundary details[28], [29], [30]. Meanwhile, Transformers [31] have significantly enhanced the modeling of long-range dependencies in remote sensing imagery through their global self-attention mechanisms. For example, SETR [32] and Segmenter [33] reframe semantic segmentation as a sequence-to-sequence prediction task, pioneering a perspective that shifts from traditional convolutional stacking to global self-attention mechanisms. However, the quadratic complexity of self-attention mechanisms [31] renders them sensitive to training data scale and computational resources, prompting researchers to propose hybrid architectures that balance modeling capability with efficiency. Therefore, some studies, such as UNetFormer [34] and ST-UNet [25], have proposed hybrid architectures combining CNNs and Transformers to balance modeling capacity and computational efficiency. Nevertheless, Zhang et al. [35] point out that both CNN-based and Transformer-based models tend to focus on low-frequency components in images, resulting in insufficient modeling of boundary information and fine-grained semantics.
@@ II-B Boundary-Aware Semantic Segmentation
To optimize model prediction accuracy in boundary regions, current research primarily focuses on designing boundary losses, constructing boundary or semantic-aware modules, or indirectly enhancing boundary representation capabilities from the frequency domain perspective. Regarding boundary losses, InverseForm [36] constructs a loss function based on boundary distance that leverages inverse transformation networks to capture spatial transformations between predicted and ground truth boundaries. Wang et al. [37] introduce a boundary loss that predicts differentiable direction vectors to guide predicted boundaries toward progressive alignment with ground truth boundaries. Wang et al. [9] employ a boundary loss function based on geometric similarity that seamlessly integrates boundary guidance into multi-level feature fusion to enhance boundary accuracy. Additionally, other research focuses on constructing boundary or semantic-aware modules to boost boundary feature discrimination. For example, Mask2Anomaly [38] reformulates the segmentation paradigm as mask-level inference. GPINet [22] constructs a geometric prior generation module with iterative updates. BFP [8] and DNLNet [39] explicitly model boundaries as independent semantic entities, while LANet [40] and HSDN [41] propose semantic decoupling strategies to separate features from different categories. These works improve semantic consistency in homogeneous and boundary regions. Furthermore, some studies approach the problem from a frequency domain perspective. For instance, FsaNet [35] proposes a linear self-attention mechanism targeting low-frequency components, significantly reducing computational overhead while enhancing intra-class consistency and edge contour information. FDNet [42] introduces frequency attention mechanisms to process high and low-frequency information separately, effectively improving image edge preservation and category discrimination capabilities. Meanwhile, PointRend [6] treats image segmentation as a graphics rendering problem and uses an iterative subdivision algorithm to adaptively select points in uncertain boundary regions, enabling point-level segmentation prediction.
However, discriminative learning models inherently tend to prioritize minimizing overall classification errors while neglecting high-frequency boundary information, resulting in insufficient learning and misclassification of boundary pixels [12]. Existing methods essentially improve the inherent limitations of discriminative models without breaking through the framework constraints of discriminative modeling.
@@ II-C Diffusion-Based Generative Models
In recent years, diffusion models have garnered widespread attention due to their powerful image generation capabilities. Some research has begun to incorporate diffusion models into segmentation tasks [43],[44],[45]. For example, Baranchuk et al. [46] discovered that intermediate activations in the reverse diffusion process of image generation can effectively capture pixel-level semantic representations for semantic segmentation. [47],[48] explored the implementation of diffusion models in segmentation tasks from a binary segmentation perspective. Additionally, studies such as [49], [50], and [51] have discovered implicitly aligned image-text representations in pretrained diffusion models and explored their applications in semantic segmentation tasks. However, from the perspective of overall segmentation metrics, there remains a significant gap in semantic reasoning accuracy between diffusion generative models and discriminative models [20]. Addressing this issue, DDP [52] replaced the traditional L2 loss in diffusion with cross-entropy loss and constructed a lightweight noise-to-prediction conditional diffusion-aware pipeline. Furthermore, Chen et al. [53] achieved competitive results by introducing bit diffusion models in panoptic segmentation tasks, formulating label generation as a discrete data generation problem.
We observe that discriminative learning excels at capturing low-frequency, large-scale semantic features but struggles to accurately learn high-frequency boundary features. Conversely, diffusion generative learning can effectively capture high-frequency boundary features but lacks accuracy in semantic feature inference. We combines the advantages of both discriminative learning and diffusion generative learning by implementing a conditional guidance network that learns joint guidance representations from coarse segmentation maps predicted by discriminative models and remote sensing images. This effectively guides the diffusion generative model to simultaneously focus on both global semantic structures and boundary details, significantly enhancing both the semantic integrity and boundary accuracy of segmentation results.

### v2

@@ Abstract
Remote sensing semantic segmentation must address both what the ground objects are within an image and where they are located. Consequently, segmentation models must ensure not only the semantic correctness of large-scale patches (low-frequency information) but also the precise localization of boundaries between patches (high-frequency information related to boundary components). However, most existing approaches rely heavily on discriminative learning, which excels at capturing low-frequency features, while overlooking its inherent limitations in learning high-frequency features for semantic segmentation. Recent studies have revealed that diffusion generative models excel at generating high-frequency details. Our theoretical analysis confirms that the diffusion denoising process significantly enhances the model’s ability to learn high-frequency features; however, we also observe that these models exhibit insufficient semantic inference for low-frequency features when guided solely by the original image. Therefore, we integrate the strengths of both discriminative and generative learning, proposing the Integration of Discriminative and diffusion-based Generative learning for Boundary Refinement (IDGBR) framework. The framework first generates a coarse segmentation map using a discriminative backbone model. This map and the original image are fed into a conditioning guidance network to jointly learn a guidance representation subsequently leveraged by an iterative denoising diffusion process refining the coarse segmentation. Extensive experiments across five remote sensing semantic segmentation datasets (binary and multi-class segmentation) confirm our framework’s capability of consistent boundary refinement for coarse results from diverse discriminative architectures.
@@ I Introduction
Semantic segmentation aims to classify each pixel in an image, thereby partitioning the image into multiple regions with specific semantic meanings. In remote sensing, semantic segmentation is also known as land use or land cover classification, representing a fundamental task in remote sensing image analysis. Semantic segmentation of remote sensing images provides crucial support for numerous remote sensing applications, including geographic information updates, environmental monitoring, and urban planning.
Figure 1: A comparative evaluation between a diffusion model and a discriminative model for semantic segmentation of remote sensing imagery was conducted, with both models fully trained on the Potsdam dataset. The comparison is carried out from two perspectives: (a) visualization from a frequency-based viewpoint, and (b) overall semantic reasoning accuracy. The results reveal complementary characteristics between the two models: the diffusion model exhibits an inherent advantage in generating fine-grained boundaries (high-frequency information), while the discriminative model demonstrates stronger capability in semantic reasoning (low-frequency information).
@@
In recent years, with the development of deep learning technology, semantic segmentation methods utilizing deep learning for remote sensing images have overcome the performance bottlenecks of traditional pixel-oriented or object-oriented segmentation approaches [1, 2]. Although these methods have significantly improved segmentation performance, they often fail to produce satisfactory boundary segmentation results [3]. Most studies attribute the difficulty of boundary region segmentation to the fact that segmentation models rely on large-scale features to capture contextual information from spatial neighborhoods [4, 5]. Compared to local features, large-scale features indeed provide richer semantic information. However, large-scale features are typically obtained from low-resolution feature maps, large receptive fields, long-range pixel relationships, or global pixel associations, thereby sacrificing local detail information. To address this issue, many studies have attempted to combine contextual semantic feature learning with boundary-sensitive local feature learning. Bertasius et al. [6] enhanced the model’s boundary localization awareness by jointly optimizing pixel-wise affinity loss and semantic segmentation loss. Ding et al. [7] introduced an additional boundary category to be optimized along with semantic categories, and utilized a directed acyclic graph structure to use the learned boundaries to isolate contextual information propagation between different regions. Wang et al. [8] designed a boundary-guided multi-scale feature fusion module to enhance boundary information preservation in multi-scale features. Considering the variability of boundary pixels between different categories, Wu et al. [9] pulled each boundary pixel closer to its nearest local category center while pushing it away from surrounding categories.
@@
However, existing studies primarily focus on enhancing intra-class similarity and inter-class separability in boundary regions under discriminative classification strategies, without adequately addressing the negative effects of discriminative learning on boundary segmentation. Boundaries refer to local transitions between adjacent different regions in the spatial domain. From a signal frequency perspective, boundaries are represented as high-frequency components in the frequency domain. Studies have shown that Convolutional Neural Networks (CNNs) can capture high-frequency components in image features [10]; however, discriminative classification tends to make CNNs prioritize learning low-frequency components while suppressing high-frequency ones [11]. This leads to insufficient feature learning for boundary pixels, thereby increasing the likelihood of their misclassification. From a probability theory perspective, discriminative classification directly learns the posterior distribution, disregarding the inherent characteristics of the data itself and only focusing on minimizing overall classification errors in the pixel space for given input data [12]. Due to the small proportion of boundary pixels in the pixel space, errors in boundary pixel classification are often overlooked or even suppressed during the discriminative learning process [13].
@@
Unlike discriminative learning, generative learning aims to produce high-fidelity samples by modeling the joint distribution of input data [12]. Therefore, theoretically, generative learning needs to accurately model both low-frequency content structures and high-frequency boundary details in images, without neglecting or suppressing the learning of high-frequency components as discriminative learning does. However, for generative models such as VAEs and GANs, learning the distribution of high-frequency components remains challenging [14]. In recent years, diffusion probabilistic models have achieved breakthrough progress in generative tasks by gradually adding noise to data and learning the reverse denoising process, thereby generating high-quality samples [15, 16, 17]. In image generation tasks, studies have found that diffusion generative models enhance the learning of high-frequency components in images [18]. This paper also observes similar experimental phenomena in semantic segmentation tasks: by using remote sensing images as conditional information and applying diffusion noise training with Stable Diffusion [15] to segmentation label map encoded features, more realistic segmentation boundaries can be obtained. As shown in Fig. 1 (a), visualizing the high-frequency structural details extracted via Fourier transform reveals that diffusion models generate boundaries closer to the target geometric contours compared to discriminative models. However, from the perspective of overall segmentation metrics, there remains a significant gap in semantic reasoning accuracy between the segmentation results of diffusion generative models and discriminative models [19], as shown in Fig. 1 (b).
@@
In summary, discriminative learning excels at capturing low-frequency large-scale semantic features but struggles to accurately learn high-frequency features; conversely, diffusion generative learning can effectively capture high-frequency features but lacks accuracy in inferring semantic features. The high-frequency information in images consists of multiple components, such as noise and object boundaries. In contrast, manually annotated semantic segmentation maps are generally free from image noise, and their high-frequency information is primarily concentrated in semantic boundary regions. Based on this analysis and observation, we propose the Integration of Discriminative and diffusion-based Generative learning for Boundary Refinement (IDGBR) framework, which learns the annotation patterns in boundary regions from manually labeled segmentation maps. The framework comprises a diffusion generative model, a discriminative model, and a conditional guidance network. First, to ensure semantic accuracy of the overall segmentation, IDGBR uses coarse segmentation results predicted by a discriminative model as semantic guidance conditions for diffusion generative learning. Although these coarse segmentation results have inaccurate boundary regions, they provide reliable semantic information for interior regions. Second, IDGBR captures joint representations from coarse segmentation maps and remote sensing images via a conditional guidance network. These representations are subsequently integrated into the diffusion model through residual connections to enhance semantic and boundary fidelity. Finally, we introduce a simple regularization technique that leverages guidance from a pre-trained representation model to explicitly align intermediate generative features. This technique is designed to mitigate initial training instability and to promote the intrinsic coherence of the semantic feature space.
Notably, current mainstream semantic segmentation evaluation metrics (e.g., mIoU and F1-score) struggle to effectively assess the precision and structural integrity of predictions in boundary regions. Their fundamental limitation lies in the reliance on pixel-level statistics, which neither account for spatial dependencies between pixels nor distinguish the varying importance of prediction errors between interior and boundary regions. Consequently, traditional metrics often fail to intuitively reflect substantive improvements in repairing boundary discontinuities, correcting boundary displacements, and sharpening blurred transitions. To overcome this evaluation limitation, we introduce the boundary-sensitive evaluation metric, Weighted F-measure (WFm) [20]. By incorporating a spatially sensitive weighting mechanism, this metric enables a more accurate and targeted quantification of performance gains in terms of boundary fidelity.
The main contributions of this paper are as follows:
@@
  • We introduce a regularization technique to explicitly align diffusion generative representations with joint guidance representations, enhancing the stability and semantic coherence of the feature space during initial training.
@@ II-A General Semantic Segmentation
Remote sensing semantic segmentation often faces challenges such as insufficient boundary refinement [21, 22, 23], semantic ambiguity [1], and missed segmentation of small-scale targets [5, 24] due to large intra-class differences, low inter-class discrimination, and class imbalance in remote sensing features, as well as the large size, significant scale variations, and high scene complexity of remote sensing images. To address these issues, researchers have conducted a series of explorations.
CNNs represent the classical architecture in remote sensing semantic segmentation, with numerous research efforts building upon their strong local feature extraction capabilities to enhance the perception of multi-scale contextual information and boundary details [2, 25]. Meanwhile, Transformers [26] have significantly enhanced the modeling of long-range dependencies in remote sensing imagery through their global self-attention mechanisms. For example, SETR [27] and Segmenter [28] reframe semantic segmentation as a sequence-to-sequence prediction task, pioneering a perspective that shifts from traditional convolutional stacking to global self-attention mechanisms. However, the quadratic complexity of self-attention mechanisms [26] renders them sensitive to training data scale and computational resources, prompting researchers to propose hybrid architectures that balance modeling capability with efficiency. Therefore, some studies, such as UNetFormer [29] and ST-UNet [1], have proposed hybrid architectures combining CNNs and Transformers to balance modeling capacity and computational efficiency. Nevertheless, Zhang et al. [30] point out that both CNN-based and Transformer-based models tend to focus on low-frequency components in images, resulting in insufficient modeling of boundary information and fine-grained semantics.
@@ II-B Boundary-Aware Semantic Segmentation
To optimize model prediction accuracy in boundary regions, current research primarily focuses on designing boundary loss functions and improving model architectures. Regarding the design of boundary losses, typical approaches include geometry-aware guidance [8], distance-based inverse optimization [31], and progressive alignment via differentiable vectors [32]. However, these methods are subject to two inherent limitations [32]. First, the optimization of boundary loss is local and posterior in nature. It typically operates only within the immediate vicinity of the model’s current predictions, essentially acting as a local refinement of existing outputs. Second, boundary loss fundamentally conflicts with discriminative feature learning: while it aims to recover high-frequency details, discriminative models intrinsically favor smooth features. This contradiction leads to gradient conflicts between boundary and region losses, which often suppress boundary-related signals and result in training instability.
Regarding architectural improvements, methods can be generally categorized into two types: explicit boundary structure modeling and implicit boundary perception enhancement. The fundamental idea of explicit boundary structure modeling is to introduce a dedicated boundary processing branch within the model architecture to directly model and optimize boundary regions. These methods typically construct dual-stream network architectures, fusing conventional semantic features with boundary features through mechanisms such as gating mechanisms [33], feature propagation barrier constraints [7], or pairwise semantic disentanglement [34], thereby explicitly enhancing the model’s perception of boundary structures. Conversely, the basic idea of implicit boundary perception enhancement is not to rely on parallel geometric structure branches, but to improve feature representation learning mechanisms, enabling the model to spontaneously learn more accurate boundary representations. These methods employ strategies such as local attention embedding [35], graphics rendering concepts [5], high-order affinity-based semantic decoupling [36], task reformulation based on masked attention [37, 38], geometric prior generation [22], or frequency-aware feature enhancement [30, 39] to implicitly enhance the learning capability of boundary regions in the feature space. However, the spectral bias and texture bias [40] inherent in discriminative models are difficult to eliminate through architectural design alone. The fundamental reason lies in the optimization objective of discriminative learning, which seeks to minimize the overall error. As a result, low-frequency regions—which occupy most of the image area—dominate the gradient updates, forming an implicit inductive bias that favors low-frequency components. This bias drives the model to smooth sparse high-frequency boundaries, severely impairing the perception of global geometric topology. Consequently, architectural improvements based on discriminative models remain largely confined to “neighborhood reasoning” in terms of boundary enhancement performance.
@@ II-C Diffusion-Based Generative Models
In recent years, diffusion models have garnered widespread attention due to their powerful image generation capabilities. Some research has begun to incorporate diffusion models into segmentation tasks. For example, Baranchuk et al. [41] discovered that intermediate activations in the reverse diffusion process of image generation can effectively capture pixel-level semantic representations for semantic segmentation. Zhao et al. [42] have discovered implicitly aligned image-text representations in pretrained diffusion models and explored their applications in semantic segmentation tasks. However, from the perspective of overall segmentation metrics, there remains a significant gap in semantic reasoning accuracy between diffusion generative models and discriminative models [19]. Addressing this issue, DDP [43] replaced the traditional L2 loss in diffusion with cross-entropy loss and constructed a lightweight noise-to-prediction conditional diffusion-aware pipeline.
Similarly, in the remote sensing domain, diffusion models have demonstrated remarkable potential. In the context of remote sensing image generation, diffusion models have been successfully applied to high-fidelity remote sensing image synthesis under controllable conditions [44], continuous generation of global multi-resolution remote sensing imagery [45], and cross-modal remote sensing image editing [46]. As for remote sensing image enhancement, Ref-Diff [47] and EDiffSR [48] utilize ground object priors to explicitly guide the diffusion process, significantly improving the restoration quality of high-frequency details in super-resolution reconstruction. In summary, the successful application of diffusion models in the field of remote sensing stems from their powerful ability to learn high-frequency details, leading to superior performance over traditional methods in remote sensing image generation tasks.
We observe that discriminative learning excels at capturing low-frequency, large-scale semantic features but struggles to accurately learn high-frequency boundary features. Conversely, diffusion generative learning can effectively capture high-frequency boundary features but lacks accuracy in semantic feature inference. We combine the advantages of both discriminative learning and diffusion generative learning by implementing a conditional guidance network that learns joint guidance representations from coarse segmentation maps predicted by discriminative models and remote sensing images. This effectively guides the diffusion generative model to simultaneously focus on both global semantic structures and boundary details, significantly enhancing both the semantic integrity and boundary accuracy of segmentation results.